OpenStudy (anonymous):

Write the standard form of the equation y-3=-2/3(x+5). NEED HELP. HELP ME PLEASE!! I'm begging..

OpenStudy (texaschic101):

y - 3 = -2/3(x + 5) -- distribute through the parenthesis y - 3 = -2/3x - 10/3 -- add 3 to both sides y = -2/3x - 10/3 + 3 y = -2/3x - 10/3 + 9/3 y = -2/3x - 1/3 now change to Ax + By = C y = -2/3x - 1/3 -- add 2/3 to each side 2/3x + y = -1/3 -- multiply the entire equation by 3 2x + 3y = -1 <-- standard form any questions ?

OpenStudy (anonymous):

No, you explained that PERFECTLY!! I see where i messed up! OMG, You have like saved me! THANK YOU!!

OpenStudy (texaschic101):

no problem...glad I could help :)